Dear Richard dear all,


please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were:


Arcetri: A. Puglisi, P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
LBTO: G. Brusa
MG: M. Andrighettoni

Total # of person: 10

W unit

1) The cleaning of the WFS optical elements has been completed.
2) cleaning of the pupil rotator prism. Some additional work has been done in cleaning the accessible face of the prism contained in the pupil rotator. The spots identified as generated from this surface did not disappeared after cleaning. A more careful inspection showed that the prism coating is defective in a a few points. These points do not reflect light and generates some spots in the pupil images. The spot dimension is between one and two pixel in the pupil images and the approximate real dimension is about 0.2mm. To do a first estimate of the impact of these defects on the WFS operation a few sets of signals have been acquired with the internal reference source. The signals has been multiplied by the wavefront reconstructor measured in Arcetri. The reconstructed wavefronts have an rms in between 20/30 nm.
This is an overestimate of the total generated effect because this rms included low order modes that will not be affected by these well localized defects. This first analysis shows that the AO system performance should not be affected by these defects. These defects seems to be related to a deterioration of the prism coating that in the long term, could require replacing or re coating the prism. To start monitoring the coating conditions we will take other two set of pictures one before installing the unit on the telescope and another one before the IRTC is bolted on the AGW at the telescope focal station. See the attached ppt presentation for pictures and details.
3) The WFS cover installation is started and is about 50% completed.
4) the procedure to attach the IRTC camera to the AGW unit in the lab has been discussed with J. Little. the operation has been found possible with the present hw and has been planned for the 5th.

AO control SW

1) Switched IDL license file to laptop version because license server is down or unreachable.
2) New WFS low-level status gui, BoardGui ?, commited to svn
3) Previous days work on rerotator etc. tracking by WfsArbitrator ? commited to svn
4) Fixed multithreading bugs in Arbitrator GUIs that were causing occasional crashes.
5) Current svn version: 24. Intended to work with both AO2 and AO3 TCS builds, rotator tracking functionality on WFS enabled when working with AO3.

LBT672a

1) miniskirt removed to check the outer-ring act#255(dsp numebring): one of the two elastic contacts of the capsens board (the one compressed because the clocking error between refplate and coldplate) has the soldering broken.
2) inspection of the contact soldering of all the actuators of the outer ring: 3 actuators were found with one of the two contact with almost-broken soldering (acts#7, 567, 467, dsp numbering). The faulty elastic contacts were removed.
3) the fixed actuators have now just one elastic contact loosing redundancy. The left contact has been shaped to have larger preload assuring the contact.
4) the elevation slewing test have been repeated to check the fixed actuators, no longer problems found on those actuators
5) the elevation slewing test hes also repeated with the shell in rest position: one actuator (#147dsp) was found to touch the corresponding magnet glued on the shell when the unit is horizon pointing. The problem was fixed sliding back the actuator and increasing the magnet-to-coil gap.
6) repeating the slewing test no other problems have been found

Important remark: most of the problem found in the last days during the slewing tests and also problem summarized in the Roberto's report about non-working actuators after shipment were triggered by the reduction of clearance between actuator body and reference plate due to the possible increase of clocking error between cold-plate and reference-body. We suggest to consider some stable fix of the problem during the summer shutdown in case the operations during next months will show a progressing degradation process.

Other

A. Riccardi, G. Brusa & S. Esposito participated the LBT commissioning meeting where the plan for the AO system installation at the telescope has been reviewed.

Next update tomorrow

All the best

Simone

  • Rerotator coating:
  • rerotator-coating.jpg

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback